# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ Acknowledgements.
We thank Sheng-Han Zhou and Yi-Min Zhang for their helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, 11835009, 12422502, 12105344, 12235019, 12447101, 12073088, 11821505, 11991052, and 11947302), the China Manned Space Program (Grant Nos. CMS-CSST-2025-A01 and CMS-CSST-2025-A02), the National 111 Project (Grant No. B16009), and the National Key Research and Development Program of China (Grant Nos. 2021YFC2203004, 2021YFA0718304, and 2020YFC2201501).

### v2

@@ Acknowledgements.
We thank Sheng-Han Zhou and Yi-Min Zhang for their helpful discussions. This work was supported by the National SKA Program of China (Grants No. 2022SKA0110200 and No. 2022SKA0110203), the National Natural Science Foundation of China (Grants No. 12533001, No. 12575049, No. 12473001, No. 11975072, No. 11875102, No. 11835009, No. 12422502, No. 12105344, No. 12235019, No. 12447101, No. 12073088, No. 11821505, No. 11991052, and No. 11947302), the China Manned Space Program (Grants No. CMS-CSST-2025-A01 and No. CMS-CSST-2025-A02), the National 111 Project (Grant No. B16009), and the National Key Research and Development Program of China (Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501).
DATA AVAILABILITY
The data that support the finding of this article are openly available Ling et al. .